# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v13

@@ 2.9 Statistical Bounding
$\kappa_{s}=\sqrt{3}$ and $\kappa_{s}=5$ are defined as the ideal bounding ranges for the uniform and Gaussian distributions respectively.
@@ 9.5 Trigonometric Library Errors
For both sine functions, the uncertainty of each $\sin(x)$ is assumed to equal its ULP, which is displayed in Figure 34. Because the Quart sine function has minimal Taylor expansion error due to its minimal range of $x$ in calls to the library $\sin(x)$, its true numerical errors are likely smaller than ULP, therefore its uncertainty is sightly overestimated. In contrast, Figure 34 indicates that the Library sine function exhibits underestimated uncertainties.
@@ 9.8 Using Library Sine for Sin/Cos Signals
  • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions. This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 38 and $4$.
@@ 9.9 Numerical Error Resonance Using Library Sine
In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 48. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. For example, Figure LABEL:fig:_FFT_Cost_Clean_6_6_Wave_Lib shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 43. In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order. In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 47. Also, it is completely absent when using the Quart sine function, as seen in Figure 36 and 37. Figure 48 and Figure 49 demonstrate that the error deviations increase with sine or cosine frequency. Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 42, 43, and 46. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
@@ 9.10 Using Quart Sine for Linear Signal
Figure 50 shows that using the Quart sine function, the result uncertainties can track the value errors of Linear signals with proper coverage. Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases more rapidly with FFT order than its counterpart in Figure 38. Such increase results in larger error deviations: At FFT order $18$, $0.89>0.65$ for the forward transformation, $1.08>0.90$ for the reverse transformation, $1.67>0.90$ for the roundtrip transformation, and wider distributions of the normalized value errors in Figure 51 versus the corresponding histograms in Figure 39. Using the Quart sine function, variance arithmetic can provide proper coverage for Sin signals, Cos signals, and Linear signal.
@@ 9.12 Ideal Coverage
The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise. For Linear signals computed using the Library sine function, Figure 56 and 57 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively. Ideal coverage corresponds to the region where the error deviation equals $1$. Outside this region, proper coverage cannot be achieved for the reverse transformation. Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region. Furthermore, as numerical errors increase with computational load, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms, and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function. Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
In contrast, for Linear signals computed using the Quart sine functions, Figure 58 and 59 show the results for the forward and reverse transformations, respectively. The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions. The forward transformation exhibits a broader ideal coverage region than the reverse transformations. While the forward transformation shows consistent over-estimation of the value errors in the proper coverage region, the reverse transformation demonstrates both over-estimation and under-estimation of the value errors in the proper coverage region.
As a comparison, because Sin/Cos signals have fewer numerical calculation errors, using either Quart or the Library sine functions, the ideal coverage region is achieved once the added noise is large enough to cover the effect of rounding errors, and this condition is almost independent of FFT orders. The key difference is that when using the Quart sine functions, the error deviations within the proper coverage region differ from $1$ only marginally.
Adding noises has another obvious benefit: Different implementations of the variance arithmetic have almost identical result when the input noise is larger than $10^{-16}$. For example, Figure 58 and 59 are almost identical using either C++, or Java, or Python.
@@ 9.13 Input Uncertainty Recovery
With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 38, 40, 44, and 50, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 52. With input uncertainty unspecified, roundtrip transformation can always recover input uncertainty statistically.
@@
9.14 Prec Sine Function
Figure 63: Error deviations (left y-axis) and uncertainty mean (right y-axis) of Linear signals versus FFT orders (x-axis) and transformation types (legend) computed using the the Prec sine function.
Figure 64: Histograms of normalized errorsfor forward, reverse and roundtrip transformations (legend) of Linear signals computed using the the Prec sine function.
Figure 65: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for reverse transformations of Linear signals computed using the the Prec sine function.
The Quart sine function slightly overestimates the uncertainty, while the Library sine function underestimates it significantly. This raises the question of whether a sine function with a more accurate uncertainty quantification can be developed. If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the difference between each 128-bit value and its corresponding 64-bit value can be used as the uncertainty for the 64-bit representation. The resulting indexed sine function is referred to as the Prec sine function.
Figure 62 compares the value errors and the uncertainties of $\sin(x)^{2}+\cos(x)^{2}-1$ between the Prec sine function and the Quart sine function, and demonstrates that the Prec sine functions has smaller uncertainties numerically than the Quart sine functions, but larger value errors statically, presumably due to rounding errors introduced by the casting. This difference indicates that casting a 128-bit floating-point value to a 64-bit floating-point value is less accurate in the 64-bit floating-point arithmetic, and the Prec sine function is expected to under-perform the Quart sine function due to under-estimation of the value errors by the uncertainties.
For Sin/Cos signals, Figure 63 is very similar Figure 38, provided that the error deviations using Prec sine functions are about $1.7$ times larger than their counterparts using the Quart sine functions.
For Linear signals, when comparing Figure 64 and 51 for the normalized value error distributions:
  • For the forward transformations, using either Prec or Quart sine function has very similar distributions.
  • For the reverse transformations, using Prec sine function results in wider distributions that using Quart sine function.
Although the ideal coverage regions for the reverse transformations remains identical in Figure 65 when compared with Figure 59, using Prec sine function results in larger error deviations in the proper coverage regions. Thus, using Prec sine function is slightly inferior than using Quart sine function. In another word, casting values from higher floating-point representation to a lower floating-point arithmetic is not necessarily superior.
@@ 9.15 Calculation Response
At first, it is surprising that in Figure 66 the Java implementation is 3 times faster than the C++ implementation which has been fully optimized for speed. Because FFT contains a large amount of independent and logically identical branches, the dynamic just-in-time Java compiler [37] can optimize according to data in real-time, compile the bytecodes into native machine codes, and parallelize the execution to the multiple cores and GPUs on a modern computer. Perhaps it is necessary to rethink multi-threading as a way to manually increase the time performance, because the dynamic just-in-time optimization of Java is superior to the static optimization of C++ for repeated and heavy computations with preload data such as FFT.
@@ 11.3 Improvements Needed
The choice of ideal bounding range needs further study. For example, the choice of $\kappa_{s}=5$ for Normal distribution is based on the $5-\sigma$ rule in practice only. Such choice certainly can not be applied to other distributions.

### v14

@@ 2.9 Statistical Bounding
$\kappa_{s}=\sqrt{3}$ and $\kappa_{s}=5$ are defined as the ideal bounding ranges $\hat{\kappa}$ for the uniform and Gaussian distributions respectively.
@@ 9.5 Trigonometric Library Errors
9.6 Refine Quart Sine Function
Figure 37: The value error histogram for the Quart sine function when it is calculated using 128-bit floating-point arithmetic and cast to 64-bit floating-point representation.
If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the resulting indexed sine function is referred to as the Prec sine function.
  • Figure 36 compares the value errors and the uncertainties of $\sin(x)^{2}+\cos(x)^{2}-1$ between the Prec sine function and the Quart sine function, and demonstrates that the Prec sine functions has smaller uncertainties numerically than the Quart sine functions, but larger value errors statically, presumably due to rounding errors introduced by the casting. As suggested in Figure 36, the Prec sine function has about $5\%$ more non-zero $\sin(x)^{2}+\cos(x)^{2}-1$ value errors, and $14\%$ non-zero $\sin(x)^{2}+\cos(x)^{2}-1$ value errors exceeding the corresponding uncertainties. These difference indicate that casting a 128-bit floating-point value to a 64-bit floating-point value is less accurate in the 64-bit floating-point arithmetic.
  • Figure 37 demonstrates that the value errors of the Prec sine function after normalized by the corresponding value ULP is uniform distributed with in the half bit of ULP. Because variance arithmetic assumes that the value error is uniformly distributed within ULP of each value, the Quart sine function overestimates value errors by one times.
@@ 9.9 Using Library Sine for Sin/Cos Signals
  • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions. This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 40 and 46.
@@ 9.10 Numerical Error Resonance Using Library Sine
In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 50. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. For example, Figure LABEL:fig:_FFT_Cost_Clean_6_6_Wave_Lib shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 45. In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order. In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 49. Also, it is completely absent when using the Quart sine function, as seen in Figure 38 and 39. Figure 50 and Figure 51 demonstrate that the error deviations increase with sine or cosine frequency. Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 44, 45, and 48. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
@@ 9.11 Using Quart Sine for Linear Signal
Figure 52 shows that using the Quart sine function, the result uncertainties can track the value errors of Linear signals with proper coverage. Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases more rapidly with FFT order than its counterpart in Figure 40. Such increase results in larger error deviations: At FFT order $18$, $0.89>0.65$ for the forward transformation, $1.08>0.90$ for the reverse transformation, $1.67>0.90$ for the roundtrip transformation, and wider distributions of the normalized errors in Figure 53 versus the corresponding histograms in Figure 41. Using the Quart sine function, variance arithmetic can provide proper coverage for Sin signals, Cos signals, and Linear signal.
@@ 9.13 Ideal Coverage
The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise. For Linear signals computed using the Library sine function, Figure 58 and 59 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively. Ideal coverage corresponds to the region where the error deviation equals $1$. Outside this region, proper coverage cannot be achieved for the reverse transformation. Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region. Furthermore, as numerical errors increase with computational load, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms, and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function. Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
In contrast, for Linear signals computed using the Quart sine functions, Figure 60 and 61 show the results for the forward and reverse transformations, respectively. The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions. The forward transformation exhibits a broader ideal coverage region than the reverse transformations. While the forward transformation shows consistent over-estimation of the value errors in the proper coverage region, the reverse transformation demonstrates both over-estimation and under-estimation of the value errors in the proper coverage region. Figure 61 shows that the rounding errors balances the over-estimation of value errors in the Quart sine function in Figure 37, so that the error deviations within the proper coverage region differ from $1$ only marginally.
Adding noises has another obvious benefit: Different implementations of the variance arithmetic have almost identical result when the input noise is larger than $10^{-16}$. For example, Figure 60 and 61 are almost identical using either C++, or Java, or Python.
@@ 9.14 Input Uncertainty Recovery
With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 40, 42, 46, and 52, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 54. With input uncertainty unspecified, roundtrip transformation can always recover input uncertainty statistically.
@@ 9.15 Calculation Response
At first, it is surprising that in Figure 64 the Java implementation is 2 times faster than the C++ implementation which has been fully optimized for speed. Because FFT contains a large amount of independent and logically identical branches, the dynamic just-in-time Java compiler [37] can optimize according to data in real-time, compile the bytecodes into native machine codes, and parallelize the execution to the multiple cores and GPUs on a modern computer. Perhaps it is necessary to rethink multi-threading as a way to manually increase the time performance, because the dynamic just-in-time optimization of Java is superior to the static optimization of C++ for repeated and heavy computations with preload data such as FFT.
@@ 11.3 Improvements Needed
The choice of ideal bounding range $\hat{\kappa}$ needs further study. For example, the choice of $\hat{\kappa}=5$ for Normal distribution is based on the $5-\sigma$ rule in practice only. Such choice certainly can not be applied to other distributions. Also, the choice of ideal bounding range $\hat{\kappa}$ determines the convergence of both $\log(x)$ and $x^{c}$ as $P(x)\lesssim 1/\hat{\kappa}$, which needs better mathematical understanding.